The demand curve is QD=500-1/2P


Calculate the (point) price elasticity of demand when price is Php100.


1
Expert's answer
2021-11-04T13:48:30-0400
"E=-\frac{1}{2}\times\frac{100}{500-0.5\times100}=-\frac{1}{9}"